OpenStudy (anonymous):

simplify (x)(x+1)(x+2)=480+(x)+(x+1)+(x+2) step by step please

OpenStudy (anonymous):

\[(x)(x+1)(x+2)\text{ = }480+(x)+(x+1)+(x+2) \]\[x (1+x) (2+x)=483+3 x\]\[x^3+3 x^2+2 x-(483+3 x)=0\]\[x^3+3 x^2-x-483=0 \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

lol thanks for ur help

OpenStudy (anonymous):

You're welcome.